# Faster estimation of the transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

### v2

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. fiksel2022 proposed a transformation-free linear regression model, that minimizes the Kullback-Leibler divergence from the observed to the fitted compositions, where the EM algorithm is used to estimate the regression coefficients. We formulate the model as a constrained logistic regression, in the spirit of tsagris2025, and we estimate the regression coefficients using constrained iteratively reweighted least squares. The simulation studies depict that this algorithm makes the estimation procedure significantly faster, uses less memory, and in some cases gives a better solution.
@@ 1 Introduction
An issue with the EM algorithm adopted for the estimation of the TFLR’s coefficients is that the EM is in general slow. To mitigate this computational cost we employ the constrained iteratively reweighted least squares (CIRLS) algorithm, with simplicial constraints imposed on the regression coefficients. Tsagris’s implementation of the EM algorithm (tsagris2025) was shown to be 4 times faster than the implementation of fiksel2022. Our simulation studies demonstrate that CIRLS yields a 2-255 times, depending on the case, speed improvement over the implementation of tsagris2025. These speed improvements are particularly important in the following cases: a) having to analyze multiple datasets, b) performing simulation studies, c) working with high dimensional compositions (rios2025), and d) performing permutation or bootstrap based hypothesis tests (fiksel2022, tsagris2025).
@@ 2 The TFLR model
fiksel2022 developed the TFLR model, a novel approach for modeling compositional data, addressing key limitations in existing methodologies.
@@
Unlike conventional approaches that rely on log-ratio transformations (ALR, CLR, ILR), the TFLR model works directly in the simplex space, avoiding transformation-induced interpretability issues. The estimation method, described below, is based on first-moment specifications using estimating equations, avoiding full likelihood specifications and providing robustness to different data-generating mechanisms. The approach naturally accommodates zeros and ones in both predictors and outcomes, addressing a major limitation of log-ratio based methods that require data pre-processing. tsagris2025 though, highlighted that the aforementioned approach allows for zero values in the response variable, but not for the simplicial predictor, in general. Consider for example the case of an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are different from zero and and the $k$-th column of the estimated matrix $\bm{B}$ is equal to $(0,0,B_{3k},B_{4k})^{\top}$, where both $B_{3k}$ and $B_{4k}$ are not zero. Another possibility is when a column of the $\bm{B}$ matrix is full of zeros. These two cases will produce $x_{ik}B_{ik}=0$ and hence (4) cannot be computed. fiksel2022 bypasses this problem by adding a small quantity ($\delta=10^{-8}$) during computation of the KLD.
@@ 2.1 Expectation-maximisation algorithm
In order to avoid for() loops, by vectorization, in the EM algorithm, the implementation in the R package Compositional (compositional2025) constructs the following design matrix
$\widetilde{\bm{X}}=\bm{I}_{D_{r}}\otimes\bm{X}=\left(\begin{array}[]{cccc}\bm{X}&\bm{0}&\ldots&\bm{0}\\ \bm{0}&\ddots&\bm{0}&\vdots\\ \vdots&\bm{0}&\ddots&\bm{0}\\ \bm{0}&\ldots&\bm{0}&\bm{X}\end{array}\right).$ (9)
This resulted in an implementation that was shown to be around 4 times faster (tsagris2025) than the implementation in the R package codalm (codalm2021), at the cost of adding extra memory requirements.
@@ 3 TFLR with CIRLS
We give some preliminaries at first regarding the TFLR, the IRLS, the CIRLS and how the latter can be used to estimate the regression coefficients of the TFLR.
@@ 3.2 Logistic regression with identity link and simplicial constraints
This gives equal weight to each coefficient. Alternatively, the starting values may come from fitting a constrained OLS with simplicial constraints imposed on the regression coefficients. The convergence criterion we employed was the absolute difference in two successive values of the log-likelihood. In some cases, the optimal solution may lie on the boundary of the simplex, with some coefficients exactly equal to 0 or 1. This can lead to degeneracy in the optimization problem. QP solvers are generally designed to handle such cases, but it may be beneficial to use a solver that employs an interior point method or another approach that handles degeneracy well.
3.3 Customizing CIRLS for TFLR
In order to formulate the TFLR as a univariate logistic regression model we follow tsagris2025 and apply the same approach. They constructed the $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@ 3.3 Customizing CIRLS for TFLR
$\bm{d}=\mathop{\operator@font vec}\nolimits\left(\bm{X}^{\top}\bm{W}\cdot\bm{Y}\right),$
@@
Effectively, we do not vectorise the simplicial responses, or the compositional predictors, and evidently neither the weighting matrix $\bm{W}$. We obtain the $\bm{B}$ matrix from the QP, and subsequently we compute the matrix with the fitted responses, and then the KLD value. We then update the weighting matrix and this process iterates until convergence of the KLD. The initial matrix $\bm{B}$ is obtained from the constrained least squares method of tsagris2025. The dual active-set method of goldfarb1983, implemented in the R package quadprog (quadprog2019), is used during the QP minimization.
@@ 4.3 Implementation details in R
Table 1 presents the theoretical complexities of the two algorithms. However, the R implementations are slightly different. which is of dimensions $\left(n\times D_{rp}\right)\times\left(n\times D_{rp}\right)$. As stated earlier, the EM algorithm’s R implementation in the package Compositional (compositional2025), constructs the design matrix $\widetilde{\bm{X}}$ of Eq. (9). This save running time but increases the memory requirements to $\mathcal{O}(nD_{p}^{2}D_{r})$. As for CIRLS algorithm, its bottleneck is the computation of the block matrix $\bm{D}$ (17) that involves performing $2\times D_{r}$ cross-product matrix multiplications.
Second, computational optimizations have been applied. Instead of computing the full KLD, at each iteration of either algorithm, only the parts involving the fitted values are computed, i.e. the right hand side of Eq. (4) is computed. The matrix $\bm{D}$ (17), required by QP, is initialised once, and its elements are updated at every iteration.
@@ 5 Simulation studies
EM-1 refers to the implementation in the package Compositional, while EM-2 refers to the implementation of (fiksel2022) in the package codalm.
@@ 5.1 Speed-up factors
Figure 1 visualizes the speed-up factors of CIRLS to EM when the two compositions are independent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 1.3 up to 50, whereas compared to EM-2 they range from 8 up to 50. When $p=10$, the EM-1 is 1.6 up 220 times slower, whereas EM-2 is 12-141 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 17-50 times slower than CIRLS when $p=15$, and 22-64 times slower than CIRLS when $p=20$. Regarding the KLD, both algorithms returned the same value.
Figure 2 visualizes the speed-up factors when the compositions are dependent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 4 up to 227, whereas compared to EM-2 they range from 2 up to 96. When $p=10$, the EM-1 is 9 up 347 times slower, whereas EM-2 is 12-264 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 46-114 times slower than CIRLS when $p=15$, and 40-157 times slower than CIRLS when $p=20$. Notably, there was a discrpeancy in the produced KLD values. CIRLS always produced slightly higher KLD than the EM algorithm. The discrepancy was always in the 4rth, the 5th or the 6th digit.
@@
(e) EM-2 to CIRLS when $p=15$ (f) EM-2 to CIRLS when $p=20$
Figure 1: Speed-up factors of the two algorithms, when the two compositions are independent: running time of EM divided by the running time of CIRLS with IRLS. The horizontal golden line indicates when the unity speed-up factor, when the running times are equal.
@@
(e) EM-2 to CIRLS when $p=15$ (f) EM-2 to CIRLS when $p=20$
Figure 2: Speed-up factors of the two algorithms, when the two compositions are dependent: running time of EM divided by the running time of CIRLS with IRLS. The horizontal golden line indicates when the unity speed-up factor, when the running times are equal.
@@ 5.2 Scalability of the EM and CIRLS algorithms
To further examine and compare the scalability of the two algorithms, we estimated their scalability rate by fitting the following model $t=\alpha n^{\beta}$, or $\log(t)=\alpha+\beta n$ (goldsmith2007), and report the $\beta$ coefficient, where $t$ and $n$ denote the running time and sample size. Tables 2 and 3 present the empirically estimated scalability rates of both algorithms. Notably, all implementations exhibit linear or sublinear scalability rates, depending on the dimensionalities of the compositions.
@@
$p=5$ EM-1 0.90 0.87 0.89 0.89
EM-2 0.21 0.28 0.40 0.42
CIRLS 0.81 1.14 0.77 1.16
$p=10$ EM-1 0.79 0.72 0.76 0.73
EM-2 0.13 0.63 0.68 0.73
CIRLS 0.96 0.81 1.07 1.11
$p=15$ EM-2 0.66 0.48 0.36 0.30
CIRLS 0.82 0.70 0.60 0.54
$p=20$ EM-2 0.32 0.14 0.11 0.17
CIRLS 0.66 0.52 0.45 0.36
@@
$p=5$ EM-1 0.92 0.89 0.76 0.77
EM-2 1.01 0.91 1.03 0.99
CIRLS 0.64 0.98 0.98 0.87
$p=10$ EM-1 0.82 0.77 0.69 0.63
EM-2 1.03 1.01 0.95 0.87
CIRLS 0.88 0.98 0.97 0.90
$p=15$ EM-2 0.38 0.52 0.48 0.55
CIRLS 0.77 0.77 0.62 0.47
$p=20$ EM-2 0.73 0.66 0.67 0.64
CIRLS 0.78 0.62 0.44 0.26
@@ 6 Conclusions
We adopted the IRLS methodology used in generalised linear models to the case of the TFLR model for simplicial-simplicial regression. For this we exploited the fact that TFLR may be seen as a logistic regression model with simplicial constraints on the regression coefficients. Regardless of the dimensiolnalities in both the responses and the predictors, this resulted in significant time improvements ranging from 1.3 up to more than 300 times. The estimated scalability rates were linear or sublinear for both algorithms.
The impressive speed-up factors of CIRLS though come with a cost. When the compositions are dependent, the CIRLS fails by a small amount ($\approx 10^{-4}$) to reach the minimum KLD value. Our suggested compromise is to perform the CIRLS and use the estimated regression coefficients as initial values in the EM algorithm. For bootstrap or permutation hypothesis testing though, the CIRLS might be a good alternative.

### v3

# Scalable approximation of the transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. fiksel2022 proposed a transformation-free linear regression model, that minimizes the Kullback-Leibler divergence from the observed to the fitted compositions, where the EM algorithm is used to estimate the regression coefficients. We formulate the model as a constrained logistic regression, in the spirit of tsagris2025, and we estimate the regression coefficients using constrained iteratively reweighted least squares. The simulation studies depict that this algorithm makes the estimation procedure significantly faster, and approximates accurately enough the solution of the EM algorithm.
@@ 1 Introduction
An issue with the EM algorithm adopted for the estimation of the TFLR’s coefficients is that the EM is in generally computationally slow. To mitigate this computational cost we employ the constrained iteratively reweighted least squares (CIRLS) algorithm, with simplicial constraints imposed on the regression coefficients. Tsagris’s implementation of the EM algorithm (tsagris2025) was shown to be 4 times faster than the implementation of fiksel2022. Our simulation studies demonstrate that CIRLS yields a 2-255 times, depending on the case, speed improvement over the implementation of tsagris2025. The obtained solution is an accurate approximation of the solution produced by the EM algortihm. The speed improvements are particularly important in the following cases: a) having to analyze multiple datasets, b) performing simulation studies, c) working with high dimensional compositions (rios2025), and d) performing permutation or bootstrap based hypothesis tests (fiksel2022, tsagris2025).
@@ 2 The TFLR model
fiksel2022 developed the TFLR model, a novel approach for modeling compositional data, addressing methodological limitations inherent in existing methodologies.
@@
Unlike conventional approaches that rely on log-ratio transformations (ALR, CLR, ILR), the TFLR model works directly in the simplex space, avoiding transformation-induced interpretability issues. The estimation method, described below, is based on first-moment specifications using estimating equations, avoiding full likelihood specifications and providing robustness to different data-generating mechanisms. The approach naturally accommodates zeros and ones in both predictors and outcomes, addressing a major limitation of log-ratio based methods that require data pre-processing. However, tsagris2025 highlighted that the aforementioned approach allows for zero values in the response variable, but not for the simplicial predictor, in general. Consider for example the case of an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are different from zero and and the $k$-th column of the estimated matrix $\bm{B}$ is equal to $(0,0,B_{3k},B_{4k})^{\top}$, where both $B_{3k}$ and $B_{4k}$ are not zero. Another possibility is when a column of the $\bm{B}$ matrix is full of zeros. These two cases will produce $x_{ik}B_{ik}=0$ and hence (4) cannot be computed. fiksel2022 circumvents this limitation by adding a small quantity ($\delta=10^{-8}$) during computation of the KLD.
@@ 2.1 Expectation-maximisation algorithm
Further, the initial values in the Initialization Step are provided by the simplicially constrained least squares model of tsagris2025. This resulted in an implementation that was shown to be around 4 times faster (tsagris2025) than the implementation in the R package codalm (codalm2021), at the cost of adding extra memory requirements.
@@ 3 TFLR with CIRLS
We first present preliminary concepts regarding the TFLR, the IRLS, the CIRLS and how the latter can be used to estimate the regression coefficients of the TFLR.
@@ 3.2 Logistic regression with identity link and simplicial constraints
This assigns equal weight to each coefficient. Alternatively, the starting values may come from fitting a constrained OLS with simplicial constraints imposed on the regression coefficients. The convergence criterion we employed was the absolute difference in two successive values of the log-likelihood. In some cases, the optimal solution may lie on the boundary of the simplex, with some coefficients exactly equal to 0 or 1. This can lead to degeneracy in the optimization problem. QP solvers are generally designed to handle such cases, but it may be advantageous to use a solver that employs an interior point method or another approach that handles degeneracy well.
3.3 CIRLS formulation of the TFLR
To formulate the TFLR as a univariate logistic regression model we follow tsagris2025 and apply the same approach. They constructed the $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@ 3.3 CIRLS formulation of the TFLR
$\bm{d}=\mathop{\operator@font vec}\nolimits\left(\bm{X}^{\top}\bm{W}\cdot\bm{Y}\right),$ (13)
@@
Effectively, we do not vectorise the simplicial responses, or the compositional predictors, and evidently neither the weighting matrix $\bm{W}$. We obtain the $\bm{B}$ matrix from the QP, and subsequently we compute the matrix with the fitted responses, and then the KLD value. We then update the weighting matrix and this process iterates until convergence of the KLD. The initial matrix $\bm{B}$ is obtained from the constrained least squares method of tsagris2025.
The matrix $\bm{D}$ (12) and the vector $\bm{d}$ (13) increase the memory requirements. However, there is no option to reduce the memory requirements and also parallelize the QP step, an important consideration for high-dimensional compositional data.
TFLR solutions may lie on the boundary of the simplex, with some coefficients attaining exactly 0 or $1$. In such cases the active-set approach of (goldfarb1983) is particularly well suited because it handles degeneracy by maintaining a working set of active constraints whose feasibility is enforced at every iteration. This property ensures that the IRLS update always remains in the simplex and that convergence is not impeded by rank deficiencies in the reduced Hessian.
We employ the dual active-set implementation of the R package quadprog (quadprog2019), which is efficient when the number of active constraints at the optimum is small—an empirically common scenario when TFLR solutions lie near the interior of the simplex. Alternative interior-point methods could be used, especially in very high-dimensional regimes where the $O(D_{p}^{3})$ factorisation cost becomes dominant, but we found the active-set method to be numerically stable and highly effective for the dimensionalities considered in this work.
@@ 4.3 Implementation details in R
Table 1 presents the theoretical complexities of the two algorithms. The bottleneck of the CIRLS is the computation of the block matrix $\bm{D}$ (12) that involves performing $2\times D_{r}$ cross-product matrix multiplications. Second, computational optimizations have been applied. Instead of computing the full KLD, at each iteration of either algorithm, only the parts involving the fitted values are computed, i.e. the right hand side of Eq. (4) is computed. The matrix $\bm{D}$ (12), required by QP, is initialised once, and its elements are updated at every iteration.
@@ 5 Simulation studies
EM-1 refers to the implementation in the package codalm, EM-2 refers to the implementation in the package Compositional. The tolerance value to terminate the EM and CIRLS algorithms was set to $10^{-8}$. The simulations were performed in a desktop with Intel Core i9-14900K CPU at 3.2GHz, 128 GB RAM and Windows 11 Pro installed.
@@ 5.1 Speed-up factors
Figure 1 visualizes the speed-up factors of CIRLS to EM when the two compositions are independent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 6 up to 99, whereas compared to EM-2 they range from 6 up to 65. When $p=10$, the EM-1 is 20–157 times slower, whereas EM-2 is 13–76 times slower. When $p=15$ EM-1 is 13–30 times slower than CIRLS, and EM-2 is 24–62 times slower. Finally, when $p=20$ EM-1 is 14–51 times slower than CIRLS, and EM-2 is 27–75 times slower. Regarding the KLD, the discrepancy between EM and CIRLS was $\approx 10^{-7}-10^{-6}$.
Figure 2 visualizes the speed-up factors when the compositions are dependent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 4 up to 227, whereas compared to EM-2 they range from 2 up to 96. When $p=10$, the EM-1 is 9 up 347 times slower, whereas EM-2 is 12-264 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 46-114 times slower than CIRLS when $p=15$, and 40-157 times slower than CIRLS when $p=20$. Notably, there was a discrepancy in the produced KLD values. CIRLS always produced slightly higher KLD than the EM algorithm, the discrepancy was $\approx 10^{-6}-10^{-4}$.
@@
(e) EM-1 to CIRLS when $p=15$ (f) EM-2 to CIRLS when $p=15$
Refer to caption Refer to caption
(g) EM-1 to CIRLS when $p=20$ (h) EM-2 to CIRLS when $p=20$
Figure 1: Speed-up factors of the two algorithms, when the two compositions are independent: running time of EM divided by the running time of CIRLS. The horizontal golden line indicates the unity speed-up factor, when the running times are equal.
@@
(e) EM-1 to CIRLS when $p=15$ (f) EM-1 to CIRLS when $p=20$
Figure 2: Speed-up factors of the two algorithms, when the two compositions are dependent: running time of EM divided by the running time of CIRLS. The horizontal golden line indicates the unity speed-up factor, when the running times are equal.
@@ 5.2 Scalability of the EM and CIRLS algorithms
To further examine and compare the scalability of the two algorithms, we estimated their scalability rate by fitting the following model $t=\alpha n^{\beta}$, or $\log(t)=\alpha+\beta\log(n)$ (goldsmith2007), and report the $\beta$ coefficient, where $t$ and $n$ denote the running time and sample size. Tables 2 and 3 present the empirically estimated scalability rates of both algorithms. Notably, all implementations exhibit linear or sublinear scalability rates, depending on the dimensionalities of the compositions.
@@
$p=5$ EM-1 0.89 0.90 0.89 0.96
EM-2 0.77 0.78 0.77 0.83
CIRLS 0.74 0.92 0.93 0.86
$p=10$ EM-1 0.76 0.78 0.82 0.76
EM-2 0.79 0.75 0.75 0.71
CIRLS 1.01 1.02 0.82 0.92
$p=15$ EM-1 0.64 0.54 0.41 0.34
EM-2 0.64 0.54 0.41 0.34
CIRLS 0.75 0.71 0.67 0.57
$p=20$ EM-1 0.30 0.18 0.14 0.13
EM-2 0.62 0.62 0.63 0.62
CIRLS 0.72 0.65 0.55 0.45
@@
$p=5$ EM-1 0.92 0.91 0.83 0.66
EM-2 0.96 0.93 0.86 0.79
CIRLS 0.74 1.01 1.01 1.01
$p=10$ EM-1 0.80 0.63 0.44 0.40
EM-2 0.81 0.71 0.66 0.66
CIRLS 0.91 0.99 0.82 0.89
$p=15$ EM-1 0.69 0.68 0.65 0.64
CIRLS 0.91 0.70 0.48 0.29
$p=20$ EM-1 0.91 0.79 0.80 0.80
CIRLS 0.88 0.63 0.40 0.18
Further, both sets of simulation studies revealed that the EM-2 is slower than EM-1. Using the CIRLS produced coefficients as initial values in the EM-2 algorithm improved its speed, but not significantly.
@@ 6 Conclusions
We adopted the IRLS methodology used in generalised linear models to the case of the TFLR model for simplicial-simplicial regression. For this we exploited the fact that TFLR may be seen as a logistic regression model with simplicial constraints on the regression coefficients. Regardless of the dimensionalities in both the responses and the predictors, this resulted in significant time improvements ranging from 1.3 up to more than 300 times. The estimated scalability rates were linear or sublinear for both algorithms. The remarkable speed-up factors of CIRLS though are accompanied by a small trade-off. When the compositions are dependent, the CIRLS fails by a small amount ($\approx 10^{-6}-10^{-4}$) to reach the minimum KLD value. This discrepancy is negligible for most practical applications, especially for permutation and bootstrap hypothesis testing.
The CIRLS creates opportunities for additional methodological developments in simplicial-simplicial regression. fiksel2022 proposed to minimize the KLD from the observed to the fitted compositions, and tsagris2025 the objective function with the $L_{2}$ and $L_{1}$ norms. $\phi$-divergence regression is another direction, that includes the symmetric KLD, and the Jensen-Shannon divergence (a symmetrized version of the KLD), among others. The CIRLS algorithm could be used to approximate the exact solutions of this type of regression, yielding fast, but approximate solutions, that could also be used as initial values to an exact algorithm.